# Conformal Killing gravity in static spherically-symmetric spacetimes

### v1

@@ I Conformal Killing gravity
As $R_{kl}$ contains second order derivatives of the metric tensor, higher orders in the field equations (2) may arise with the tensor $K_{jk}$. This does not occur in the present work, as well as in [3], where we obtained a realization of the conformal Killing parametrization in FRW background. In [3] the CKT has the perfect fluid form, candidate for representing the dark sector, and contains the scale factor with no derivatives. The Friedmann equations are thus second order in the metric, and reproduced the same forecasting obtained by Harada [2] with eqs.(1). Vacuum cosmological solutions, wormhole and black hole solutions were obtained by Clemént and Nouicer [9]. In FRW background, CKG is embraced by a general parametrization of Codazzi tensors [10].
@@
Most papers on CKG dealt with a static background. Barnes [12] found the general spherically symmetric vacuum solution and [13] the general solution with Maxwell source. Junior, Lobo and Rodriguez [14] investigated regular black holes solutions of CKG coupled to non -linear electrodynamics and scalar fields. Further in [15] they explored black bounce solutions in CKG coupled to non -linear electrodynamics and scalar fields. pp-wave solutions were studied by Barnes [16].
@@
We show that the (unique) Schwarzschild-like solution is the vacuum by Harada. and remark that it cannot originate from a perfect fluid CKT.
In section V we write the CKG equations for nonlinear electrodynamics. The same result for the metric functions in Sect. III is proven in the non-linear case.
@@ II Anisotropic CKT
$\nabla_{k}u_{j}=-u_{j}\dot{u}_{j},\qquad\nabla_{j}\dot{u}_{k}=\nabla_{k}\dot{u}_{j}$ (4)
@@
$R=R^{k}{}_{k}$ and the curvature $R^{\star}$ of the space-like submanifold are eqs. 90, 91:
@@ III.1 Properties of metric functions
$\nabla_{p}\dot{u}^{p}=\dfrac{1}{h^{2}}(bb^{\prime}+b^{\prime\prime})-\dfrac{bb^{\prime}}{h^{2}}\dfrac{h^{\prime}}{h}+\dfrac{2bb^{\prime}}{rh^{2}}$ (19)
$\Sigma=-\dfrac{1}{h^{2}}(bb^{\prime}+b^{\prime\prime})+bb^{\prime}\dfrac{h^{\prime}}{h^{3}}+\dfrac{1}{r^{2}}\dfrac{b^{2}}{h^{2}}-\dfrac{1}{r^{2}}+\dfrac{b^{2}}{r}\dfrac{h^{\prime}}{h^{3}}$ (20)
@@ V.1 Linear electrodynamics
This is eq.14 worked by Barnes in [13] and the “master equation” by Clemént and Noucier [9].
@@
As a check, the second order equation (37) is solved for $\kappa_{4}=1$, $x=r\sqrt{\kappa}_{3}$. The solution, evaluated with Mathematica, coincides with that by Clemént and Noucier [9]
@@ VI Conclusions
The CKG equations can support an anisotropic fluid source as well as (non)-linear electrodynamics. In both cases the equations are second order. This is a great advantage with respect to the solution by components of the third order Harada equations found in the existing literature. We prove the equivalence of our second order equations with the third order ones. Our approach recovers several results obtained so far, in a simple and covariant way, and give some new ones.
@@ Appendix 1. Useful formulae
$\frac{\chi^{s}\nabla_{s}\eta}{2\sqrt{\eta}}=\frac{1}{2\sqrt{\eta}}\frac{1}{f_{1}}\frac{d}{dr}\left(\frac{b^{\prime 2}}{b^{2}f_{1}^{2}}\right)$ (53)
$=\frac{1}{f_{1}^{2}}\left[\frac{b^{\prime\prime}}{b}-\frac{b^{\prime 2}}{b^{2}}-\frac{b^{\prime}f_{1}^{\prime}}{bf_{1}}\right]$
@@
$\nabla_{j}\chi_{k}=\frac{\nabla_{j}\dot{u}_{k}}{\sqrt{\eta}}-\frac{\chi^{s}\nabla_{s}\eta}{2\eta}\chi_{j}\chi_{j}$ (54)
@@
The following relations (see eqs 16 and 19 of [18] in the static spherically symmetric case) are to be considered:

### v3

@@ I Conformal Killing gravity
As $R_{kl}$ contains second order derivatives of the metric tensor, higher orders in the field equations (2) may arise with the tensor $K_{jk}$. This does not occur in the present work, as well as in [3], where we obtained a realization of the conformal Killing parametrization in FRW background. In [3] the CKT has the perfect fluid form, candidate for representing the dark sector, and contains the scale factor with no derivatives. The Friedmann equations are thus second order in the metric, and reproduced the same forecasting obtained by Harada [2] with eqs.(1). Vacuum cosmological solutions, wormhole and black hole solutions were obtained by Clément and Nouicer [9]. In FRW background, CKG is embraced by a general parametrization of Codazzi tensors [10].
@@
Most papers on CKG dealt with a static background. Barnes [12] found the general spherically symmetric vacuum solution and [13, 14] the general solution with Maxwell source. Junior, Lobo and Rodriguez [15] investigated regular black holes solutions of CKG coupled to nonlinear electrodynamics and scalar fields. Further in [16] they explored black bounce solutions in CKG coupled to nonlinear electrodynamics and scalar fields. pp-wave solutions were studied by Barnes [17].
@@
We show that the (unique) Schwarzschild-like solution is the vacuum by Harada and remark that it cannot originate from a perfect fluid CKT.
In section V we write the CKG equations for nonlinear electrodynamics. The same result for the metric functions in Sect. III is proven in the nonlinear case.
@@ II Anisotropic CKT
$\nabla_{k}u_{j}=-u_{k}\dot{u}_{j},\qquad\nabla_{j}\dot{u}_{k}=\nabla_{k}\dot{u}_{j}$ (4)
@@
$R=R^{k}{}_{k}$ and the curvature $R^{\star}$ of the space-like submanifold are (eqs. 90, 91 of [18]):
@@ III.1 Properties of metric functions
$\nabla_{p}\dot{u}^{p}=\dfrac{1}{h^{2}}(b^{\prime 2}+bb^{\prime\prime})-\dfrac{bb^{\prime}}{h^{2}}\dfrac{h^{\prime}}{h}+\dfrac{2bb^{\prime}}{rh^{2}}$ (19)
$\Sigma=-\dfrac{1}{h^{2}}(b^{\prime 2}+bb^{\prime\prime})+bb^{\prime}\dfrac{h^{\prime}}{h^{3}}+\dfrac{1}{r^{2}}\dfrac{b^{2}}{h^{2}}-\dfrac{1}{r^{2}}+\dfrac{b^{2}}{r}\dfrac{h^{\prime}}{h^{3}}$ (20)
@@ V.1 Linear electrodynamics
This is eq.14 worked by Barnes in [13] and the “master equation” by Clément and Noucier [9].
@@
As a check, the second order equation (37) is solved for $\kappa_{4}=1$, $x=r\sqrt{\kappa}_{3}$. The solution, evaluated with Mathematica, coincides with that by Clément and Noucier [9]
@@ VI Conclusions
The CKG equations can support an anisotropic fluid source as well as (non)-linear electrodynamics. In both cases the equations are second order. This is a great advantage with respect to the solution by components of the third order Harada equations found in the existing literature. We prove the equivalence of our second order equations with the third order ones. Our approach recovers several results obtained so far in a simple and covariant way, and gives some new ones.
@@ Appendix 1. Useful formulae
$\frac{\chi^{s}\nabla_{s}\eta}{2\sqrt{\eta}}=\frac{1}{2\sqrt{\eta}}\frac{1}{f_{1}}\frac{d}{dr}\left(\frac{b^{\prime 2}}{b^{2}f_{1}^{2}}\right)$
$=\frac{1}{f_{1}^{2}}\left[\frac{b^{\prime\prime}}{b}-\frac{b^{\prime 2}}{b^{2}}-\frac{b^{\prime}f_{1}^{\prime}}{bf_{1}}\right]$ (53)
@@
$\nabla_{j}\chi_{k}=\frac{\nabla_{j}\dot{u}_{k}}{\sqrt{\eta}}-\frac{\chi^{s}\nabla_{s}\eta}{2\eta}\chi_{j}\chi_{k}$ (54)
@@
Eq. 87 in [18], eqs.(7) and (53) give:
@@
Data availability statement
No data are available because of the nature of the research.